basiamorgan:

A 76.00 76.00 pound flask of mercury costs $151.00. $151.00. The density of mercury is 13.534 g/cm 3 . 13.534 g/cm3. Find the price of one cubic inch of mercury by calculating intermediate values. What is the price of one pound of mercury?

Vocaloid:

price of one pound of mercury ---> simply divide (total $ spent)/(total # of pounds), to get the price per pound now, for the harder question, the price per 1 cubic inch. thinking about this in terms of units, we want the final unit to be $/cm^3 start with the units we're given \[\frac{ $151.00 }{ 76.00~lbs }\]from there, we can go from lbs to grams. to do this, we multiply by the conversion rate (1lb/453.592g, since 1 lb = 453.592g), so we can cancel out lbs and end up with grams at the bottom. \[\frac{ $151.00 }{ 76.00~lbs }\frac{ 1~lb }{ 453.592g }\] now, we want to go from grams to cubic centimers. recall that we are also given the density of mercury 13.534 g/cm^3. this is perfect, because we want to cancel out grams, and end up with cm^3 in the bottom of our expression. \[\frac{ $151.00 }{ 76.00~lbs }\frac{ 1~lb }{ 453.592g }\frac{ 13.534 g }{ cm^3 }\] simplify this expression, cancel out units, and you should end up with the price per cm^3

Vocaloid:

Addendum I’m an idiot, it wants it in cubic inches not cubic cm 2.54 cm = 1 inch Therefore 2.54^3 cm^3 = 1 inch^3 Take the previous value, multiply by the conversion factor 2.54^3 cm^3 / 1 inch^3